OpenStudy (softballgirl372015):

Can someone please help me verify this trig identity! cos^2-sin^2=1-2sin^2

OpenStudy (softballgirl372015):

Do you know how to do this problem?

OpenStudy (anonymous):

for alegbra?

OpenStudy (softballgirl372015):

I am doing preacalc

OpenStudy (anonymous):

cos^2 =1- sin^2 therefore cos^2 -sin^2 = 1 -sin^2-sin^2 =1- 2sin^2 that's it

OpenStudy (softballgirl372015):

How did you do that? When verifying trig identities you have to only simplify one side of the equal sign.

OpenStudy (softballgirl372015):

Oh. Nevermind! I get what your saying!! Thanks! :)